OpenStudy (anonymous):

Compute \[\int_0^1 \frac{x^r-1}{\ln(x)} \, dx \] Where \(r>0\) is a real number.

OpenStudy (anonymous):

Level Calculus III

ganeshie8 (ganeshie8):

I think DUI/feynman trick will do here \[\begin{align} I(r)&=\int_0^1 \frac{x^r-1}{\ln(x)} \, dx\\~\\ I'(r)&=\int_0^1 x^r \, dx = \dfrac{1}{r+1}\\~\\ I(r) &= \ln(r+1) + C \end{align}\] \(I(0) = 0 \implies C = 0\)

OpenStudy (kainui):

I dare someone to try to find a method that doesn't involve differentiation under the integral sign. =P

OpenStudy (kainui):

(I am not claiming that another method exists, but I only suspect there might be, just for fun)

OpenStudy (kainui):

Well I am really technically doing the same thing as ganeshie I believe, but I'll go ahead and post it anyways. Originally I was looking to try to do some thing with the geometric series but since r is any real number that really won't work. At any rate, here we go: \[\Large x^r-1 = x^r-x^0 = x^y |_0^r = \int\limits_0^rx^y \ln x dy\] Replace this in the integral as \[\Large \int\limits_0^1 \int\limits_0^r x^y dydx = \int\limits_0^r \int\limits_0^1 x^ydxdy \\ \Large \int\limits_0^r \frac{1}{y+1}dy\] Yeah, just sort of ended up looking at it inside out oh well lol.

OpenStudy (anonymous):

By a change of variable, this can be transformed to integral that I posted in my previous problem.

OpenStudy (anonymous):

\[ x= e^u\\ dx= e^u du \\ \int_0^1 \frac{x^r-1}{\ln(x)} \, dx= \int_0^\infty \frac{e^{-(r+1)u}- e^{-u} }{u} \, du =\ln(r+1) \]

ganeshie8 (ganeshie8):

Ahh I see this problem is a nice application of frullani's thm http://math.stackexchange.com/questions/61828/proof-of-frullanis-theorem